OpenStudy (anonymous):

Solve for x. Express the answer in simplest form. 7/1 = 3/4x -1

OpenStudy (anonymous):

how can i do this?

pooja195 (pooja195):

Hi there! Welcome to openstudy :D 7/1 = 3/4x -1 7/1 is really 7 7 = 3/4x -1 now do you have any ideas?

OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

or 1x?

OpenStudy (anonymous):

im bad at this stuff

pooja195 (pooja195):

What did you do to get 3x?

OpenStudy (anonymous):

i think i messed up...

pooja195 (pooja195):



pooja195 (pooja195):

7+1=?

OpenStudy (anonymous):

8

pooja195 (pooja195):

good now we have..